OpenStudy (anonymous):

− 29π/6 Can someone help me find the reference angle for this problem.

OpenStudy (dumbcow):

easier way... change it into mixed number \[-\frac{29\pi}{6} = -4 \frac{5}{6} \pi\] 4 pi is 2 times around circle = -5pi/6 

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

yw